OpenStudy (anonymous):

Is the coordinate (1, 2) a solution to the system? y = x + 1 2x - y = 0

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

just plug in

OpenStudy (anonymous):

2=1+1 2(1)-2=0

OpenStudy (anonymous):

dunno lets check \[2=1+1\checkmark \\2\times 1-2=0\checkmark \]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

np:)